# PANGAEA: A Global and Inclusive Benchmark for Geospatial Foundation Models

### v1

@@ 2.2 Geospatial Foundation Models
As part of the SSL learning paradigm, various studies have examined self-distillation based approaches, like in the DINO [78], where two different views are fed into two encoders, known as the student and the teacher, and the student learns to predict the teacher’s output. For example, DINO-MM [113] takes concatenated SAR-optical images as raw inputs and feeds the transformed augmented views to the student-teacher network. X-STARS [114] adds a cross-sensor alignment objective based on the DINO framework, for both pre-training from scratch and continual pre-training. There are also works [6, 115, 116, 117, 118, 119] contributing to the importance of different pre-training steps. For instance, MTP [6] conducts a multi-task pre-training pipeline that supports both convolutional neural networks and vision transformer foundation models. DeCUR [120] decouples representations, distinguishing inter and intra-modal embeddings based on Barlow Twins [117], while RSMamba dense [118] and RSMamba [119] are two approaches based on Mamba [121], using selective state spaces. [122] surveys models and trends in vision-language modeling, where RemoteCLIP [123] was one of the first vision-language foundation models to learn aligned text embeddings and visual features for remote sensing. In addition, generative foundation models have also advanced within the EO field, examples include DiffusionSAT [124], CRS-Diff [125] and MetaEarth [126], which generate realistic EO images all over the Earth based on the diffusion models.
@@ 2.3 Benchmarks in Geospatial Tasks
Several benchmarks also evaluate vision-language models (VLMs) for EO applications. For example, VRSBench [146] datasets set a benchmark for image captioning, visual grounding, and visual question answering using remote sensing images with paired captions, object references, and question-answers. UrBench [147] evaluates multi-modal capabilities in urban environments, with questions spanning geo-localization, scene understanding, and object understanding. VLEO-Bench [148] presents 11 real-world scenarios to test VLMs’ abilities in scene understanding, localization and counting, and change detection. However, evaluating VLMs falls outside the scope of this work. Our benchmark focuses on dense-label tasks crucial for fine-grained spatial understanding, whereas vision-language tasks often center around image captioning, high-level scene interpretation, or object recognition, which differ from the dense prediction tasks we emphasize. In the next section, we outline the parameters we have selected to evaluate the significance of a benchmark and describe the methodology we employed to identify the most effective one.
@@ 4 Experimental Protocol
agraphHyperparameters For a fair comparison across models, the decoder is fine-tuned with an identical set of hyperparameters. For every model and dataset combination, the decoder is fine-tuned for 80 epochs using the Adam optimizer with an initial learning rate of 1e-4, a batch size of 8, and a weight decay of 0.05. The learning is decreased by a factor of 0.1 using a multi-step schedule at 60% and 90% of the training. Some specific edge cases are detailed in Appendix A for runs that necessitate adjustments due to memory constraints. The best-performing checkpoint on the validation set is kept and used to evaluate the model on the test set. For evaluation, to process larger images, we employ a sliding window inference strategy, dividing the input image into evenly distributed smaller crops. A stride is used to ensure the crops collectively cover the entire image with the fewest patches, and the overlapping areas are averaged.
@@ 5.1 Main Comparison
Model HLS Burns MADOS PASTIS Sen1Floods11 xView2 FBP DynEarthNet CropMap SN7 AI4Farms BioMassters #Top2 CROMA 82.42 67.55 32.32 90.89 39.78 51.83 38.29 49.38 59.28 25.65 36.81 2 DOFA 80.63 59.58 30.02 89.37 51.86 43.18 39.29 51.33 61.84 27.07 42.81 2 GFM-Swin 76.90 64.71 21.24 72.60 49.48 67.18 34.09 46.98 60.89 27.19 46.83 1 Prithvi 83.62 49.98 33.93 90.37 30.74 46.81 27.86 43.07 56.54 26.86 39.99 1 RemoteCLIP 76.59 60.00 18.23 74.26 48.85 69.19 31.78 52.05 57.76 25.12 49.79 2 SatlasNet 79.96 55.86 17.51 90.30 39.22 50.97 36.31 46.97 61.88 25.13 41.67 0 Scale-MAE 76.68 57.32 24.55 74.13 51.33 67.19 35.11 25.42 62.96 21.47 47.15 3 SpectralGPT 80.47 57.99 35.44 89.07 30.75 33.42 37.85 46.95 58.86 26.75 36.11 1 S12-MoCo 81.58 51.76 34.49 89.26 36.06 53.02 35.44 48.58 57.64 25.38 40.21 0 S12-DINO 81.72 49.37 36.18 88.61 33.68 51.15 34.81 48.66 56.47 25.62 41.23 1 S12-MAE 81.91 49.90 32.03 87.79 33.51 51.92 34.08 45.8 57.13 24.69 41.07 0 S12-Data2Vec 81.91 44.36 34.32 88.15 33.93 48.82 35.90 54.03 58.23 24.23 41.91 1 UNet Baseline 84.51 54.79 31.60 91.42 46.23 60.47 39.46 47.57 62.09 46.34 35.67 6 ViT Baseline 81.58 48.19 38.53 87.66 39.12 59.32 36.83 44.08 52.57 38.37 38.55 2
@@
Limitations in Highly Imbalanced and Complex Datasets. GFMs –and the supervised baselines– struggle in tasks with a high class imbalance and complex segmentation requirements. For example, on the xView2 dataset, all the models eventually collapse to predicting one or two out of five classes, indicating that even with high model capacity, performance is strongly impacted by dataset imbalance. We had to make changes to our default training setup to obtain model checkpoints that produce non-trivial outputs (see Section A.5). But even with these changes, we are still more than 25% mIoU below the state-of-the-art results [164] for this dataset. This suggests that GFMs may struggle to generalize to heavily skewed tasks like xView2, which requires nuanced multi-object segmentation and more custom approaches to achieve high performance. This limitation underscores the need for strategies specifically targeting class imbalance, as GFMs, still alleviating drawbacks like domain shift, alone may not suffice for complex and unbalanced tasks.
@@ 5.2 Data Scarcity
Model HLS Burns MADOS PASTIS Sen1Floods11 xView2 FBP DynEarthNet CropMap SN7 AI4Farms BioMassters # Top-2 CROMA 81.52 57.68 32.33 90.57 39.54 48.01 38.30 42.20 59.31 28.19 38.50 4 DOFA 78.02 55.21 28.60 88.39 50.91 36.90 39.20 30.93 47.06 26.69 42.81 2 GFM-Swin 74.36 63.37 20.41 71.61 49.06 63.14 31.25 31.42 59.83 28.43 48.19 2 Prithvi 80.89 40.79 33.13 89.69 29.88 40.27 33.43 42.51 49.45 29.27 41.03 1 RemoteCLIP 74.28 53.26 17.46 71.67 49.67 65.92 30.91 36.3 50.83 25.11 50.09 1 SatlasNet 75.97 52.24 16.78 89.45 40.94 46.04 36.34 35.29 60.74 27.08 42.23 1 Scale-MAE 75.47 46.87 23.26 72.54 53.29 62.11 32.60 20.32 61.24 26.40 46.74 2 SpectralGPT 76.40 58.00 34.61 87.52 29.50 21.71 36.52 32.09 56.28 27.46 37.34 2 S12-MoCo 79.79 42.90 32.59 89.22 32.09 46.92 34.45 41.32 56.21 28.38 41.08 0 S12-DINO 80.12 40.42 35.71 88.93 31.98 44.85 32.76 31.13 55.14 25.68 41.47 1 S12-MAE 80.13 44.29 31.15 88.43 31.00 45.63 33.29 28.07 55.55 27.50 41.66 0 S12-Data2Vec 79.82 41.22 33.42 86.58 32.46 46.73 32.61 28.53 56.94 25.84 42.82 0 UNet Baseline 82.39 43.87 30.25 90.91 49.56 55.42 35.14 36.30 46.82 45.02 36.72 4 ViT Baseline 78.17 28.77 38.71 86.08 40.40 57.32 37.33 39.53 49.21 38.37 39.56 2
@@
Model HLS Burns MADOS PASTIS Sen1Floods11 xView2 FBP DynEarthNet CropMap SN7 AI4Farms BioMassters #Top2 CROMA 76.44 32.44 32.80 87.22 29.88 37.39 36.08 36.77 42.15 38.48 40.25 6 DOFA 71.98 23.77 27.68 82.84 50.10 27.82 39.15 29.91 46.10 27.74 46.03 2 GFM-Swin 67.23 28.19 21.47 62.57 48.36 55.58 28.16 27.21 39.48 32.88 49.30 0 Prithvi 77.73 21.24 33.56 86.28 23.50 29.98 32.28 27.71 36.78 35.04 41.19 0 RemoteCLIP 69.40 20.57 17.19 62.22 48.21 56.23 34.43 19.86 43.11 23.85 53.32 1 SatlasNet 74.79 29.87 16.76 83.92 29.15 37.86 34.64 29.08 49.78 13.91 44.38 2 Scale-MAE 75.47 21.47 22.86 64.74 51.62 48.75 35.27 13.44 49.68 26.66 54.16 2 SpectralGPT 83.35 20.29 34.53 83.12 25.26 39.51 35.33 31.06 36.31 37.35 39.44 2 S12-MoCo 73.11 19.47 32.51 79.58 26.31 35.57 32.24 36.54 49.46 37.97 44.83 0 S12-DINO 75.93 23.47 36.62 84.95 27.50 34.63 32.78 38.44 41.15 37.91 42.74 2 S12-MAE 76.60 18.44 31.06 84.81 27.75 35.56 30.59 35.29 40.51 23.60 43.76 0 S12-Data2Vec 74.38 17.86 33.09 81.91 27.23 37.27 33.63 34.11 40.66 22.85 46.52 0 UNet Baseline 79.46 24.30 29.53 88.55 29.93 52.58 35.59 13.88 46.08 34.84 40.39 2 ViT Baseline 75.92 10.18 38.44 81.85 39.94 56.53 35.39 27.76 36.01 39.20 44.89 3
@@ 5.8 Impact of Encoder Fine-tuning
xView2 DOFA 51.86 50.15 -3% Prithvi 30.74 40.55 +32%
@@ A.4.1 Extensive Results
Figure A6: Example data of Five Billion Pixels (Location: China. Domain: Urban Land Cover).
@@ A.5.1 Extensive Results
Details on experimental setup: To counteract the strong class imbalance, we follow a simple oversampling approach from the winning solution [165] of the xView2 competition: Images without buildings are seen once per training epoch, images with undamaged buildings are seen twice, and images with damaged buildings are seen thrice. Furthermore, when taking random crops in training, we consider multiple candidates and prefer those with more pixels of the underrepresented classes. As the decoder model, we use a UPerNet that concatenates the two embeddings computed for the pre- and post-disaster images by the encoder. For the UNet baseline, we use an input side length of 256 pixels. The ViT uses an input side length of 224 pixels. Since the models all tended to collapse to only predicting the background class in preliminary experiments, we deviate slightly from the default setup: Instead of evaluating the model on the validation set every five epochs, we validate after every single epoch, to find a useful model checkpoint before the model collapses into a local minimum from which it does not escape anymore. Furthermore, we increase the batch size from the default 8 to 32, to decrease the likelihood of batches that contain no or very few buildings, and thus lead the model into the undesired local minimum. For the 100% experiments, increasing the batch size leads to an average increase in mIoU of 3.43% among non-baseline GFM models, with a maximum of 7.79% for CROMA and a minimum of -0.15% for Prithvi.
@@
CROMA 39.78 51.16 95.17 43.61 96.14 10.26 21.63 27.27
DOFA 51.86 64.15 97.05 61.90 97.76 16.52 36.32 46.80
GFMSwin 49.48 61.68 96.84 60.10 97.47 14.86 32.75 42.24
Prithvi 30.74 39.32 94.80 26.05 95.59 6.41 12.06 13.58
RemoteCLIP 48.85 61.13 96.62 58.09 97.33 14.68 32.84 41.29
SatlasNet 39.22 50.31 95.48 43.73 96.57 10.71 17.90 27.18
Scale-MAE 51.33 63.89 97.11 61.77 97.87 19.54 34.89 42.60
SpectralGPT 30.75 38.98 94.53 27.94 95.26 5.79 5.48 19.28
S12-MoCo 36.06 46.28 95.25 39.48 95.96 7.43 18.99 18.44
S12-DINO 33.68 43.03 94.35 37.62 95.36 7.25 16.00 12.18
S12-MAE 33.51 42.89 94.51 35.76 95.54 8.05 17.35 10.85
S12-Data2Vec 33.93 43.03 94.90 40.16 95.61 5.48 15.35 13.04
UNet Baseline 46.23 58.20 96.57 54.64 97.54 12.73 27.40 38.85
ViT Baseline 53.09 65.53 97.08 61.96 97.79 18.04 39.95 47.74
The full results in Table A7 for 100% of the data show that especially the minor and major damage classes are very hard to classify, which is also why we are oversampling them during training. No model achieves $\geq$20IoU on minor damage, which is likely due to the dataset imbalance and the difficulty of distinguishing this class from no damage or major damage. Scale-MAE performs consistently strongly, similar to DOFA, GFM-Swin, and RemoteCLIP. All of them are pre-trained on high-resolution data, which makes them well-equipped to extract features from the high-resolution xView2 images. On the other hand, SpectralGPT, the S12 models and Prithvi perform less well, likely due to them being pre-trained on lower-resolution data. This reinforces the intuition that there should be a match between the GFM and the target dataset to achieve a good performance.
Compared to the UPerNets that are trained on top of the frozen GFM encoders, the two fully trained baselines perform surprisingly well. The UNet performs roughly as well as an average GFM model, which is surprising, given that it has fewer parameters than a UperNet decoder alone. The pre-trained ViT model, on the other hand, achieves the best mIoU across all models. While it is very similar in per-class performances to the other top models, it achieves a lead of 3.63% on the major damage class. This is likely a result of being able to fine-tune the features, thus being able to recognize the fine-grained differences that define the boundaries between minor and major damage and full destruction of a building, which the more general GFM features are not attuned to.

### v2

@@ 2.2 Geospatial Foundation Models
As part of the SSL learning paradigm, various studies have examined self-distillation based approaches, like in the DINO [78], where two different views are fed into two encoders, known as the student and the teacher, and the student learns to predict the teacher’s output. For example, DINO-MM [113] takes concatenated SAR-optical images as raw inputs and feeds the transformed augmented views to the student-teacher network. X-STARS [114] adds a cross-sensor alignment objective based on the DINO framework, for both pre-training from scratch and continual pre-training. There are also works [6, 115, 116, 117, 118, 119] contributing to the importance of different pre-training steps. For instance, MTP [6] conducts a multi-task pre-training pipeline that supports both convolutional neural networks and vision transformer foundation models. DeCUR [120] decouples representations, distinguishing inter and intra-modal embeddings based on Barlow Twins [117], while RSMamba dense [118] and RSMamba [119] are two approaches based on Mamba [121], using selective state spaces. [122] surveys models and trends in vision-language modeling, where RemoteCLIP [123] was one of the first vision-language foundation models to learn aligned text embeddings and visual features for remote sensing. Other works explored the use of large language models (LLMs) for spatio-temporal modeling and prediction tasks [124, 125]. In addition, generative foundation models have also advanced within the EO field, examples include DiffusionSAT [126], CRS-Diff [127] and MetaEarth [128], which generate realistic EO images all over the Earth based on the diffusion models.
@@ 2.3 Benchmarks in Geospatial Tasks
Several benchmarks also evaluate vision-language models (VLMs) for EO applications. For example, VRSBench [148] datasets set a benchmark for image captioning, visual grounding, and visual question answering using remote sensing images with paired captions, object references, and question-answers. UrBench [149] evaluates multi-modal capabilities in urban environments, with questions spanning geo-localization, scene understanding, and object understanding. VLEO-Bench [150] presents 11 real-world scenarios to test VLMs’ abilities in scene understanding, localization and counting, and change detection. However, evaluating VLMs falls outside the scope of this work. Our benchmark focuses on dense-label tasks crucial for fine-grained spatial understanding, whereas vision-language tasks often center around image captioning, high-level scene interpretation, or object recognition, which differ from the dense prediction tasks we emphasize.
In the next section, we outline the parameters we have selected to evaluate the significance of a benchmark and describe the methodology we employed to identify the most effective one.
@@ 4 Experimental Protocol
Hyperparameters
For a fair comparison across models, the decoder is fine-tuned with an identical set of hyperparameters. For every model and dataset combination, the decoder is fine-tuned for 80 epochs using the Adam optimizer with an initial learning rate of 1e-4, a batch size of 8, and a weight decay of 0.05. The learning is decreased by a factor of 0.1 using a multi-step schedule at 60% and 90% of the training. Some specific edge cases are detailed in Appendix A for runs that necessitate adjustments due to memory constraints. The best-performing checkpoint on the validation set is kept and used to evaluate the model on the test set. For evaluation, to process larger images, we employ a sliding window inference strategy, dividing the input image into evenly distributed smaller crops. A stride is used to ensure the crops collectively cover the entire image with the fewest patches, and the overlapping areas are averaged.
@@ 5.1 Main Comparison
Model HLS Burns MADOS PASTIS Sen1Floods11 xView2 FBP DynEarthNet CropMap SN7 AI4Farms BioMassters #Top2 CROMA 82.42 67.55 32.32 90.89 53.27 51.83 38.29 49.38 59.28 25.65 36.81 2 DOFA 80.63 59.58 30.02 89.37 59.64 43.18 39.29 51.33 61.84 27.07 42.81 2 GFM-Swin 76.90 64.71 21.24 72.60 59.15 67.18 34.09 46.98 60.89 27.19 46.83 1 Prithvi 83.62 49.98 33.93 90.37 49.35 46.81 27.86 43.07 56.54 26.86 39.99 1 RemoteCLIP 76.59 60.00 18.23 74.26 57.41 69.19 31.78 52.05 57.76 25.12 49.79 2 SatlasNet 79.96 55.86 17.51 90.30 52.23 50.97 36.31 46.97 61.88 25.13 41.67 0 Scale-MAE 76.68 57.32 24.55 74.13 60.72 67.19 35.11 25.42 62.96 21.47 47.15 3 SpectralGPT 80.47 57.99 35.44 89.07 48.40 33.42 37.85 46.95 58.86 26.75 36.11 1 S12-MoCo 81.58 51.76 34.49 89.26 51.59 53.02 35.44 48.58 57.64 25.38 40.21 0 S12-DINO 81.72 49.37 36.18 88.61 50.56 51.15 34.81 48.66 56.47 25.62 41.23 1 S12-MAE 81.91 49.90 32.03 87.79 50.44 51.92 34.08 45.8 57.13 24.69 41.07 0 S12-Data2Vec 81.91 44.36 34.32 88.15 51.36 48.82 35.90 54.03 58.23 24.23 41.91 1 UNet Baseline 84.51 54.79 31.60 91.42 58.68 60.47 39.46 47.57 62.09 46.34 35.67 6 ViT Baseline 81.58 48.19 38.53 87.66 57.43 59.32 36.83 44.08 52.57 38.37 38.55 2
@@ 5.2 Data Scarcity
Model HLS Burns MADOS PASTIS Sen1Floods11 xView2 FBP DynEarthNet CropMap SN7 AI4Farms BioMassters # Top-2 CROMA 81.52 57.68 32.33 90.57 51.44 48.01 38.30 42.20 59.31 28.19 38.50 4 DOFA 78.02 55.21 28.60 88.39 58.91 36.90 39.20 30.93 47.06 26.69 42.81 2 GFM-Swin 74.36 63.37 20.41 71.61 57.81 63.14 31.25 31.42 59.83 28.43 48.19 2 Prithvi 80.89 40.79 33.13 89.69 45.79 40.27 33.43 42.51 49.45 29.27 41.03 1 RemoteCLIP 74.28 53.26 17.46 71.67 57.43 65.92 30.91 36.3 50.83 25.11 50.09 1 SatlasNet 75.97 52.24 16.78 89.45 50.74 46.04 36.34 35.29 60.74 27.08 42.23 1 Scale-MAE 75.47 46.87 23.26 72.54 59.45 62.11 32.60 20.32 61.24 26.40 46.74 2 SpectralGPT 76.40 58.00 34.61 87.52 45.94 21.71 36.52 32.09 56.28 27.46 37.34 2 S12-MoCo 79.79 42.90 32.59 89.22 49.66 46.92 34.45 41.32 56.21 28.38 41.08 0 S12-DINO 80.12 40.42 35.71 88.93 48.46 44.85 32.76 31.13 55.14 25.68 41.47 1 S12-MAE 80.13 44.29 31.15 88.43 47.09 45.63 33.29 28.07 55.55 27.50 41.66 0 S12-Data2Vec 79.82 41.22 33.42 86.58 48.84 46.73 32.61 28.53 56.94 25.84 42.82 0 UNet Baseline 82.39 43.87 30.25 90.91 56.58 55.42 35.14 36.30 46.82 45.02 36.72 4 ViT Baseline 78.17 28.77 38.71 86.08 54.82 57.32 37.33 39.53 49.21 38.37 39.56 2
@@
Model HLS Burns MADOS PASTIS Sen1Floods11 xView2 FBP DynEarthNet CropMap SN7 AI4Farms BioMassters #Top2 CROMA 76.44 32.44 32.80 87.22 46.54 37.39 36.08 36.77 42.15 38.48 40.25 6 DOFA 71.98 23.77 27.68 82.84 55.60 27.82 39.15 29.91 46.10 27.74 46.03 2 GFM-Swin 67.23 28.19 21.47 62.57 53.45 55.58 28.16 27.21 39.48 32.88 49.30 0 Prithvi 77.73 21.24 33.56 86.28 35.08 29.98 32.28 27.71 36.78 35.04 41.19 0 RemoteCLIP 69.40 20.57 17.19 62.22 53.75 56.23 34.43 19.86 43.11 23.85 53.32 1 SatlasNet 74.79 29.87 16.76 83.92 44.07 37.86 34.64 29.08 49.78 13.91 44.38 2 Scale-MAE 75.47 21.47 22.86 64.74 56.06 48.75 35.27 13.44 49.68 26.66 54.16 2 SpectralGPT 83.35 20.29 34.53 83.12 35.81 39.51 35.33 31.06 36.31 37.35 39.44 2 S12-MoCo 73.11 19.47 32.51 79.58 41.15 35.57 32.24 36.54 49.46 37.97 44.83 0 S12-DINO 75.93 23.47 36.62 84.95 41.02 34.63 32.78 38.44 41.15 37.91 42.74 2 S12-MAE 76.60 18.44 31.06 84.81 39.84 35.56 30.59 35.29 40.51 23.60 43.76 0 S12-Data2Vec 74.38 17.86 33.09 81.91 41.60 37.27 33.63 34.11 40.66 22.85 46.52 0 UNet Baseline 79.46 24.30 29.53 88.55 46.77 52.58 35.59 13.88 46.08 34.84 40.39 2 ViT Baseline 75.92 10.18 38.44 81.85 44.85 56.53 35.39 27.76 36.01 39.20 44.89 3
@@ 5.8 Impact of Encoder Fine-tuning
xView2 ScaleMAE 60.72 62.16 +2% Prithvi 49.35 57.33 +16%
@@ A.4.1 Extensive Results
Figure A6: Example data of Five Billion Pixels (Location: China. Domain: Urban Land Cover). The label colors indicate land cover categories.
@@ A.5.1 Extensive Results
Details on experimental setup: To counteract the strong class imbalance, we follow a simple oversampling approach from the winning solution [166] of the xView2 competition: Images without buildings are seen once per training epoch, images with undamaged buildings are seen twice, and images with damaged buildings are seen thrice. Furthermore, when taking random crops in training, we consider multiple candidates and prefer those with more pixels of the underrepresented classes. As the decoder model, we use a UPerNet that concatenates the two embeddings computed for the pre- and post-disaster images by the encoder. For the UNet baseline, we use an input side length of 256 pixels. The ViT uses an input side length of 224 pixels. Due to optimization issues that occurred when using Dice loss in initial experiments, we deviate from the default training protocol when training models on xView2: Instead of evaluating the model on the validation set every five epochs, we validate after every single epoch. Furthermore, we increase the batch size from the default 8 to 32, to decrease the likelihood of batches that contain no or very few buildings, and thus lead the model into the undesired local minimum. We use the cross-entropy loss for all experiments on xView2. For Table 12, we use the second-worst model Prithvi instead of the worst model SpectralGPT, due to the large computational requirements of training the whole SpectralGPT encoder.
@@
CROMA 53.27 65.81 97.21 97.77 60.64 19.97 36.13 51.85
DOFA 59.64 71.73 97.88 98.33 70.22 25.27 45.45 58.91
GFMSwin 59.15 71.45 97.73 98.19 68.11 25.92 45.45 58.06
Prithvi 49.35 61.92 96.92 97.36 56.66 16.54 35.13 41.05
RemoteCLIP 57.41 69.87 97.58 98.01 65.75 24.05 43.76 55.50
SatlasNet 52.23 64.66 97.14 97.76 59.96 19.25 32.43 51.78
Scale-MAE 60.72 72.70 97.99 98.45 71.43 26.18 47.76 59.78
SpectralGPT 48.40 60.89 96.72 97.36 54.20 17.06 28.64 44.74
S12-MoCo 51.59 64.29 97.09 97.57 58.71 18.97 37.56 45.14
S12-DINO 50.56 63.38 96.95 97.39 57.03 19.36 36.26 42.77
S12-MAE 50.44 63.05 96.99 97.44 57.95 17.44 36.15 43.24
S12-Data2Vec 51.36 64.14 97.03 97.49 57.97 19.29 38.32 43.73
UNet Baseline 58.68 70.80 97.82 98.34 68.97 23.62 44.49 57.96
ViT Baseline 57.43 70.00 97.53 97.91 65.26 24.36 46.74 52.87
The full results in Table A7 for 100% of the data show that especially the minor and major damage classes are very hard to classify, which is also why we are oversampling them during training. No model achieves $\geq$30IoU on minor damage, which is likely due to the dataset imbalance and the difficulty of distinguishing this class from no damage or major damage. Scale-MAE performs consistently strongly, similar to DOFA, GFM-Swin, and RemoteCLIP. All of them are pre-trained on high-resolution data, which makes them well-equipped to extract features from the high-resolution xView2 images. On the other hand, SpectralGPT, the S12 models and Prithvi perform less well, likely due to them being pre-trained on lower-resolution data. This reinforces the intuition that there should be a match between the GFM and the target dataset to achieve a good performance.
Compared to the UPerNets that are trained on top of the frozen GFM encoders, the two fully trained baselines perform surprisingly well. The UNet performs roughly as well as the GFMs trained on high-resolution images, which is surprising, given that it has fewer parameters than a UperNet decoder alone. The ViT model performs similarly well, while using ten times as many parameters as the UNet.